One of the uranium 235 nuclear fission reactions in a nuclear reactor is the following: n10 + U23592 = Zr91 40 + La14258 + x n10 + y e0-1 . Determine the number of x neutrons and the number of y neutrons produced by this reaction.

1
Expert's answer
2021-03-04T11:51:30-0500

La has a number 57 is the periodic table, 58 is Ce.

"\frac{1}{0}n + \frac{235}{92}U = \frac{91}{40}Zr + \frac{142}{57}La + 3\frac{1}{0}n + 5\frac{0}{-1}e"